Two forces And B acts at a point at right angles if their resultant is 50N and their sum is 70N calculate their magnitude


1
Expert's answer
2021-03-17T16:54:37-0400
"y=70-x\\\sqrt{x^2+(70-x)^2}=50\\x=40\ N\\y=70-40=30\ N"